OpenStudy (anonymous):

G(x)= 4Sin((pi)(x))-3 ((Pi)(x))-3= 0 -> got 3/Pi for this one Don't get this one ((Pi)(x))-3=((2)(Pi)) I got 5. Is rah correct? Please help.

OpenStudy (anonymous):

\[4 \sin(\pi x )-3=0\]

OpenStudy (anonymous):

is this u wish to solve?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\sin(\pi x)=3/4\] \[\sin (\pi/3)=\sqrt{3}/2\]

OpenStudy (anonymous):

Im sorry, this wasn't what i was looking for. can you look at my other question